OpenStudy (anonymous):

For all real numbers x, the function f(x) satisfies 2f(x)+f(1-x)=x^2. Find f(5)??

OpenStudy (anonymous):

34/3

OpenStudy (anonymous):

how have you found that???

OpenStudy (anonymous):

f(1-x) = x^2 - 2f(x).

OpenStudy (anonymous):

substitute x -> 1-x

OpenStudy (anonymous):

2f(1-x) + f(x) = (1-x)^2

OpenStudy (anonymous):

2[x^2 - 2f(x)] + f(x) = (1-x)^2, ie: 2x^2 - 3f(x) = (1-x)^2, ie: 3f(x) = 2x^2 - (1-x)^2 = x^2 + 2x - 1

OpenStudy (anonymous):

f(x) = (x^2 + 2x - 1) / 3

OpenStudy (anonymous):

f(5)= ?? (u knw it well) by substituting 5

OpenStudy (anonymous):

you get 34/3

OpenStudy (anonymous):

oh ok i understand it now... thanx a lot))

OpenStudy (anonymous):

come on maths chat